OpenStudy (anonymous):

Which number is a solution to x2 + 8x + 12 = 0

OpenStudy (anonymous):

combine like terms so 10x+12=0 subtract 12 -12 -12 divide by 10 10x=-12 x= 1.2

OpenStudy (haseeb96):

x^2+2x+6x+12=0 x(x+2)+6(x+2)=0 (x+6)(x+2)=0 either x+6=0 or x+2=0 x=-6 or x =-2 this is the correct answer